OpenStudy (anonymous):

If a 3-letter "word" is formed by randomly choosing 3 letters from the word OCEAN, what is the probability that it is composed only of vowels?

OpenStudy (kropot72):

First draw: P(vowel) = 3/5 Second draw: P(vowel) = 2/4 Third draw: P(vowel) = 1/3 \[P(3\ vowels)=\frac{3\times 2\times 1}{5\times 4\times 3}=you\ can\ calculate\]